# Which Reasoning Trajectories Teach Students to Reason Better? A Simple Metric of Informative Alignment

### v1

@@ 5.2 Teacher Selection
The teacher selection task aims to identify the most suitable teacher model for generating reasoning trajectories to train a given student. We consider a low-resource setting in which generating full training data for every teacher model is infeasible. Instead, we sample a small set of 200 trajectories from each candidate teacher, score them using different metrics, and select the teacher model based on the average trajectory-level score. We use 6 diverse teacher models (Deepseek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, and GPT-OSS-20B) as the candidate pool, avoiding consistently well-performing teachers to ensure a non-trivial selection task.
@@ A.1 Determining Training Problem Set
All teacher trajectory datasets are constructed using a fixed problem set to avoid confounding effects from variations in problem composition. We focus on mathematical reasoning and curate 5,000 math problems from the widely used NuminaMath dataset [21]. Following prior work [20], we apply preprocessing steps such as difficulty filtering to ensure the problem quality.
Specifically, our training dataset consists of problems drawn from the MATH, AIME/AMC, and Olympiads. SkyThought provides a difficulty-labeled version of NuminaMath in which each problem is assigned an integer difficulty score from 0 to 9. Using this scale, we randomly sampled 1,667 problems from the MATH subset with difficulty above three, 1,667 problems from the Olympiads subset with difficulty above eight, and 1,666 problems from the AIME/AMC subset, yielding a balanced training set of 5,000 problems.
We choose 5,000 training problems for three main reasons. First, prior studies have shown that strong reasoning capabilities can be learned from training sets of around 1,000 problems. Second, high-quality reasoning problems are relatively scarce, making further scaling less effective in practice. Third, our cross-study incurs quadratic computational cost across teacher–student pairs. Considering these factors, we find 5,000 problems to offer a good trade-off between representativeness and computational feasibility.
@@ A.2 Teacher Trajectory Generation
If a sampled trajectory exceeds the token budget, we resample up to 10 times; if all attempts still exceed the budget, we truncate the final trajectory. Across all teacher models, truncation rates are below 1%, which helps preserve the quality of training trajectories even for teachers that tend to produce long outputs.
@@ A.3 Teacher Models and Student Models
As discussed earlier, our experiments adopt a more diverse set of teacher models than prior work. Specifically, we consider the teacher variants shown in Table 8.
@@ A.5 Details of Model Fine-Tuning
We use LLaMA-Factory [50] framework to conduct SFT with reasoning trajectory datasets. For different student models, we conduct grid search for best set of hyperparameters. The final setting is reported in the Table 9. During SFT, all student models use a maximum sequence length of 32,768.
Most experiments are conducted on NVIDIA H200 GPUs. A single SFT experiment using the 14B model takes approximately 6 hours on 8 H200 GPUs for training.
@@ A.7 Details of Correlation Analysis
We use the Spearman correlation coefficient [45] because our analysis focuses on monotonic consistency rather than strict linear relationships between metrics and post-training performance. We also report Pearson correlation results [7] in § C.2.
@@ B.3 Influence Score
We adopt the Influence Score method based on the second-order approximation of Influence Functions [16]. The core principle is to treat the infinitesimal up-weighting of a training sample as a local perturbation to the optimization objective, thereby estimating the induced marginal change in the loss of the evaluation set. In this framework, a higher influence score signifies that the gradient direction of the training sample aligns with minimizing the evaluation loss in the neighborhood of the converged parameters.
We implement this baseline using the Kronfluence framework. To make the computation tractable for LLMs, we employ the EK-FAC strategy to approximate the Hessian matrix [10]. Specifically, we first precompute the EK-FAC factors on a reference model. Since we do not know which trajectory is better a priori, we use the model trained with randomly selected trajectories per problem as the reference model, corresponding to the “Random” variant in Table 6. Then, we compute the pairwise influence scores between the training samples and the evaluation dataset. To optimize for memory and computational efficiency, we apply a low-rank approximation to the query gradients (rank =4=4) and utilize bfloat16 precision for the Inverse Hessian-Vector Product calculations. Finally, we average these pairwise scores across the evaluation set to obtain the sample-level influence score $s(d)$, which serves as our baseline for ranking data suitability.
@@ B.4 G-Norm
G-Norm [29] assesses the suitability of generated data for a student model by measuring the magnitude of the student’s local gradient signal induced by the data. Formally, G-Norm calculates the magnitude of the loss gradient with respect to the student model’s parameters. To address the computational constraints associated with high-dimensional parameter spaces, the method utilizes random projection for dimensionality reduction. For each reasoning trajectory $(x,y)$, we compute the gradient of the loss derived from the student model and apply length normalization to eliminate bias towards longer sequences. These gradients are then projected onto a lower-dimensional subspace via a fixed random matrix, followed by the computation of their $L_{2}$ norms.
@@ B.6 Others
Here we give formal definitions for Avg-Surprisal and Avg-Surplocal. Given a trajectory $x=(t_{1},\ldots,t_{T})$,
@@ C.1 Computational Cost for Rank-Surprisal Ratio
The computation of Rank-Surprisal Ratio requires only a single forward pass through the student model. During the forward pass, we collect each token’s surprisal and rank from the model logits. Computing Rank-Surprisal Ratio over the 5,000-trajectory dataset typically takes about one hour on a single H200 GPU, which is significantly cheaper than SFT.
@@ C.2 Additional Results for Correlation Analysis
Metrics Student Models
@@
Table 10 compares Spearman and Pearson correlations. The results show that Rank-Surprisal Ratio also exhibits strong Pearson correlation with post-training performance.
@@ C.3 Additional Ablation Study
The “Fixed student” variant computes RSR using a fixed model (Qwen-3-14B) instead of the target student, and the resulting drop in correlation highlights the importance of student-specific estimation. “Avg-Rank (clip at $r_{max}=100$)” applies rank clipping to Avg-Rank, but it still fails to achieve better correlation than rank alone. “Rank Minus Surprisal” is computed by subtracting the surprisal from the (clipped) rank. Although this metric also models the relationship between rank and surprisal, this simple subtraction does not align well with post-training performance. Rank–Entropy Ratio is computed as the ratio between the (clipped) average rank and the average token entropy, where entropy is loosely related to surprisal. While it achieves good correlation, its correlation is still lower than that of RSR.
We also experiment with different exponent choices when computing Rank-Surprisal Ratio (with the default setting corresponding to power 1 for both rank and surprisal), for example "use Rank1.05". The results indicate that our metric is robust to these variations and may yield higher correlation with hyperparameter tuning. Nevertheless, we use the default power-1 setting to keep the formulation simple.
@@ C.4.3 Analysis of Datasets Selected by RSR
Deepseek-R1 Q3-235B GPT-120B Nemotron QwQ Q3-30B Magistral GPT-20B Phi-4 Q3-8B Q3-4B RSR
RSRmin on Q3-14B 6.2% 4.5% 0.1% 0.1% 67.3% 6.7% 1.4% 0.0% 2.9% 2.1% 8.6% 2.57
RSRmin on L3.1-8B 5.5% 1.0% 0.0% 3.7% 48.8% 3.9% 9.1% 0.1% 0.2% 21.3% 6.5% 2.69
RSRmin on Q2.5-7B 4.1% 1.0% 0.7% 1.7% 55.7% 5.3% 5.9% 0.2% 0.8% 17.6% 6.9% 2.67
RSRmin on Q3-4B 2.7% 2.2% 0.0% 0.3% 76.6% 4.1% 3.1% 0.0% 0.6% 4.5% 6.1% 2.56
RSRmin on Q2.5-3B 2.4% 0.8% 0.0% 2.4% 45.1% 4.0% 12.2% 0.0% 0.2% 26.0% 6.6% 2.73
Table 14: Data composition and metrics of trajectory datasets selected by RSR in the trajectory selection experiments (§ 5.1) across different student models. Model names are abbreviated as Q for Qwen and L for LLaMA.
@@
The RSR values of the selected datasets are also reported in Table 14. These datasets consistently achieve substantially lower RSR values than the teacher trajectory datasets (see § C.2 for complete metrics), validating that our selection procedure effectively identifies trajectories with low RSR for each problem.
@@ C.4.4 Additional Trajectory Selection Experiments with Fewer Teacher Models
Table 15: Comparison of different trajectory selection methods under a reduced-teacher setting.

### v2

@@ 5.2 Teacher Selection
The teacher selection task aims to identify the most suitable teacher model for generating reasoning trajectories to train a given student model prior to distillation. We consider a realistic low-resource setting where generating full training data for every candidate teacher model is either costly or infeasible. Instead, we sample a small set of 200 trajectories from each candidate teacher, score them using different metrics, and select the teacher model based on the resulting dataset-level average score. We use 6 diverse teacher models (Deepseek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, and GPT-OSS-20B) as the candidate pool, avoiding consistently well-performing teachers to ensure a non-trivial selection task.
@@ A.1 Determining Training Problem Set
All teacher trajectory datasets are constructed using a fixed problem set to avoid confounding effects from variations in problem composition, enabling a more controlled study. We focus on mathematical reasoning and curate 5,000 math problems from the widely used NuminaMath dataset [22]. Following prior work (Sky-T1) [21], we apply preprocessing steps such as difficulty filtering to ensure the problem quality.
Specifically, our training dataset consists of problems drawn from the MATH, AIME/AMC, and Olympiads. Sky-T1 provides a difficulty-labeled version of NuminaMath in which each problem is assigned an integer difficulty score from 0 to 9. Using this scale, we randomly sampled 1,667 problems from the MATH subset with difficulty above three, 1,667 problems from the Olympiads subset with difficulty above eight, and 1,666 problems from the AIME/AMC subset, yielding a balanced training set of 5,000 problems.
We fix the training set size at 5,000 problems for three reasons. First, prior studies [45, 29] have shown that strong reasoning capabilities can be learned from training sets of around 1,000 problems. Second, high-quality reasoning problems are relatively scarce, making further scaling often unrealistic in practice. Third, our distillation study incurs quadratic computational costs across teacher–student pairs. Considering these factors, we find 5,000 problems to offer a good trade-off between representativeness and computational feasibility.
@@ A.2 Teacher Trajectory Generation
For each teacher model, we generate reasoning trajectories for 5,000 problems over three independent runs using vLLM [20] under a maximum generation budget of 31,000 tokens. Each independent run (rollout) produces a dataset of 5,000 problem-trajectory pairs, yielding 11 × 3 = 33 datasets in total.
@@
We release all 33 generated trajectory datasets on Hugging Face. [^3] https://huggingface.co/datasets/Umean/RSR_data
@@ A.3 Teacher Models and Student Models
As discussed earlier, our experiments adopt a more diverse set of teacher models than prior work. Specifically, we consider the teacher variants shown in Table 8. This set includes several recently emerged reasoning models, extending beyond the DeepSeek-R1 and QwQ teachers commonly used in prior studies.
These teachers produce trajectories that vary in length, informational content, and style. For example, trajectories generated by GPT-OSS models tend to be concise, whereas those from DeepSeek-R1 are generally more detailed. Concrete trajectory examples from different teachers can be inspected in our Hugging Face datasets referenced above.
@@ A.5 Details of Model Fine-Tuning
We perform SFT on reasoning trajectory datasets using the LLaMA-Factory [52] framework and FlashAttention-2 [8]. For different student models, we conduct grid search for best set of hyperparameters. The final setting is reported in the Table 9. During SFT, all student models use a maximum sequence length of 32,768.
Most experiments are conducted on NVIDIA H200 GPUs. A single SFT experiment using the 7B model takes around 7 hours on 8 H200 GPUs for training.
@@ A.7 Details of Correlation Analysis
We use the Spearman correlation coefficient [47] because our analysis focuses on monotonic consistency rather than strict linear relationships between metrics and post-training performance. We also report Pearson correlation [7] results in § C.2.
@@
A.8 Definition of Dataset-level RSR
For dataset-level RSR, we adopt a surprisal-weighted averaging scheme analogous to the trajectory-level weighted average (§ 3.4). This design aims to mitigate numerical instability caused by trajectories with disproportionately small average surprisal and to emphasize trajectories with larger average surprisal during dataset-level aggregation. We now formally define the dataset-level RSR used in the correlation analysis (§ 4.1, 4.2).
Let $\mathbf{X}=\{\mathbf{x}_{j}\}_{j=1}^{|\mathbf{X}|}$ denote a dataset of trajectories, where each trajectory $\mathbf{x}_{j}$ contains response tokens $\{t_{j,k}\}_{k=1}^{|\mathbf{x}_{j}|}$. For brevity, we define the trajectory-level average clipped rank and average surprisal as
$\overline{r}(\mathbf{x}_{j})\;=\;\frac{1}{|\mathbf{x}_{j}|}\sum_{k=1}^{|\mathbf{x}_{j}|}\min\!\bigl(\mathrm{Rank}(t_{j,k}),r_{max}\bigr),\qquad\overline{s}(\mathbf{x}_{j})\;=\;\frac{1}{|\mathbf{x}_{j}|}\sum_{k=1}^{|\mathbf{x}_{j}|}\mathrm{Surprisal}(t_{j,k}).$ (8)
Accordingly, the trajectory-level RSR can be equivalently written as $\mathrm{RSR}(\mathbf{x}_{j})=\frac{\overline{r}(\mathbf{x}_{j})}{\overline{s}(\mathbf{x}_{j})}$.
To obtain the dataset-level RSR, we apply a weighted averaging scheme analogous to Eq. 6, using the trajectory-level average surprisal $\overline{s}(\mathbf{x}_{j})$ as the weight:
$\mathrm{RSR}_{\text{dataset}}(\mathbf{X})$ $=\;\frac{\sum_{j=1}^{|\mathbf{X}|}\overline{s}(\mathbf{x}_{j})\,\mathrm{RSR}(\mathbf{x}_{j})}{\sum_{j=1}^{|\mathbf{X}|}\overline{s}(\mathbf{x}_{j})}$ (9)
$=\;\frac{\sum_{j}\overline{s}(\mathbf{x}_{j})\,\frac{\overline{r}(\mathbf{x}_{j})}{\overline{s}(\mathbf{x}_{j})}}{\sum_{j}\overline{s}(\mathbf{x}_{j})}$
$=\;\frac{\sum_{j}\overline{r}(\mathbf{x}_{j})}{\sum_{j}\overline{s}(\mathbf{x}_{j})}.$
The resulting dataset-level metric admits a concise form, which can be interpreted as the ratio of the summed trajectory-level average clipped ranks to the summed trajectory-level average surprisals.
We compare the correlation strength of the simple dataset-level average of trajectory-level RSR, i.e., $\frac{1}{|\mathbf{X}|}\sum_{j=1}^{|\mathbf{X}|}\mathrm{RSR}(\mathbf{x}_{j})$, with that of the dataset-level RSR computed via surprisal-weighted averaging (Eq. 9) in the additional ablation study (§ C.3).
@@ B.3 Influence Score
We adopt the Influence Score method based on the second-order approximation of influence functions [17]. The core idea is to model an infinitesimal up-weighting of a training sample as a local perturbation to the optimization objective, thereby estimating the resulting marginal change in the evaluation-set loss. Within this framework, a higher influence score indicates that the gradient direction induced by the training sample is more aligned with minimizing the evaluation loss in the vicinity of the converged parameters.
We implement this baseline using the Kronfluence framework. To make the computation tractable for LLMs, we employ the EK-FAC strategy to approximate the Hessian matrix [11]. Specifically, we first precompute the EK-FAC factors on a reference model. Since we do not know which trajectory is better a priori, we use the model trained with randomly selected trajectories per problem as the reference model, corresponding to the “Random” variant in Table 6. Then, we compute the pairwise influence scores between the training samples and the evaluation dataset. To optimize for memory and computational efficiency, we apply a low-rank approximation to the query gradients (rank =4=4) and utilize bfloat16 precision for the Inverse Hessian-Vector Product calculations. Finally, we average these pairwise scores across the evaluation set to obtain the sample-level (trajectory-level) influence score $s(d)$, which is used as the baseline for evaluating data suitability.
@@ B.4 G-Norm
G-Norm [30] assesses the suitability of generated data for a student model by measuring the magnitude of the student’s local gradient signal induced by the data. Formally, G-Norm calculates the magnitude of the loss gradient with respect to the student model’s parameters. To address the computational constraints associated with high-dimensional parameter spaces, the method utilizes random projection for dimensionality reduction. For each reasoning trajectory $\mathbf{x}$, we compute the gradient of the loss derived from the student model and apply length normalization to eliminate bias towards longer sequences. These gradients are then projected onto a lower-dimensional subspace via a fixed random matrix, followed by the computation of their $L_{2}$ norms.
@@ B.6 Others
Here we give formal definitions for Avg-Surprisal and Avg-Surplocal3.2). Given a trajectory $x=(t_{1},\ldots,t_{T})$,
@@ C.1 Computational Cost for Rank-Surprisal Ratio
The computation of Rank-Surprisal Ratio requires only a single forward pass through the student model, during which we collect each token’s surprisal and clipped rank from the model logits. For a single trajectory, computing token-level surprisals and ranks has a worst-case time complexity of $\mathcal{O}(TV)$, where $T$ denotes the number of response tokens and $V$ is the vocabulary size. This computation does not introduce additional complexity beyond that inherent in the dimensionality of the model logits. Moreover, since RSR only depends on clipped ranks, the rank computation can be efficiently implemented via top-$k$ selection rather than full-vocabulary comparisons, reducing practical runtime and GPU memory usage.
In practice, computing RSR over the 5,000-trajectory dataset with a context length of 32,768 using a 7B model typically takes under one hour on a single H200 GPU with FlashAttention-2, which is significantly cheaper than SFT.
@@ C.2 Additional Results for Correlation Analysis
Metric Correlation Measure Student Models
@@
Table 10 reports both Spearman and Pearson correlations for Rank-Surprisal Ratio. The results show that RSR also exhibits strong Pearson correlation with post-training performance. The rationale for primarily using Spearman correlation rather than Pearson correlation is discussed in § A.7.
@@ C.3 Additional Ablation Study
The “Fixed student” variant computes RSR using a fixed model (Qwen-3-14B) instead of the target student, and the resulting drop in correlation highlights the importance of student-specific estimation. The "Simple dataset-level average" variant computes the dataset-level RSR by simply averaging trajectory-level scores, as noted in § A.8. The slight degradation in correlation indicates that a simple average of trajectory-level RSR remains a robust estimator of dataset-level suitability, while the surprisal-weighted averaging scheme improves the reliability of the aggregated metric at both trajectory-level and dataset-level.
The “Avg-Rank (clipped)” variant applies rank clipping with $r_{max}=100$ to Avg-Rank; however, it still exhibits a notable performance gap compared with RSR. This suggests that rank information alone is insufficient to yield strong correlation, underscoring the importance of jointly leveraging rank and surprisal. The “Rank Minus Surprisal” metric is computed by subtracting the surprisal from the (clipped) rank. Although it also models the relationship between rank and surprisal, this simple subtraction does not align well with post-training performance. "Rank–Entropy Ratio" is defined as the ratio between the average (clipped) token rank and the average token entropy, where entropy is loosely related to surprisal. While this metric achieves relatively strong correlation, its performance remains below that of RSR.
We also experiment with different exponent choices when computing Rank-Surprisal Ratio (with the default setting corresponding to a power of 1 for both rank and surprisal), for example "use Rank1.05". The results indicate that our metric is robust to these variations and may yield higher correlation with hyperparameter tuning. Nevertheless, we use the default power-1 setting to keep the formulation simple.
@@ C.4.3 Analysis of Datasets Selected by RSR
R1 Q3-235B GPT-120B Nemotron QwQ Q3-30B Magistral GPT-20B Phi-4 Q3-8B Q3-4B RSR Length
RSRmin on Q3-14B 6.2% 4.5% 0.1% 0.1% 67.3% 6.7% 1.4% 0.0% 2.9% 2.1% 8.6% 2.57 9363
RSRmin on L3.1-8B 5.5% 1.0% 0.0% 3.7% 48.8% 3.9% 9.1% 0.1% 0.2% 21.3% 6.5% 2.69 9939
RSRmin on Q2.5-7B 4.1% 1.0% 0.7% 1.7% 55.7% 5.3% 5.9% 0.2% 0.8% 17.6% 6.9% 2.67 9845
RSRmin on Q3-4B 2.7% 2.2% 0.0% 0.3% 76.6% 4.1% 3.1% 0.0% 0.6% 4.5% 6.1% 2.56 9419
RSRmin on Q2.5-3B 2.4% 0.8% 0.0% 2.4% 45.1% 4.0% 12.2% 0.0% 0.2% 26.0% 6.6% 2.73 10169
Table 14: Data composition and metric statistics of trajectory datasets selected by RSR in the trajectory selection experiments (§ 5.1) for different student models. Model names are abbreviated as Q for Qwen and L for LLaMA.
@@
The average RSR values of the selected datasets are also reported in Table 14. These datasets consistently achieve substantially lower RSR values than the teacher trajectory datasets (see § C.2 for complete metrics), validating that our selection procedure effectively identifies trajectories with low RSR for each problem.
@@ C.4.4 Additional Trajectory Selection Experiments with Fewer Teacher Models
Table 15: Comparison of different trajectory selection methods under a reduced-teacher setting. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.